1. If 10.0 grams of an unknown metal undergoes a temperature change of 8.5°C and absorbs 32.47 J. Calculate the specific heat of the metal and identify the metal.
Answer #1

We know that,

Q = w*C*\DeltaT

Specific heat (C) = Q/(w*\DeltaT)

Given that,

Q = 32.47 J

w = 10 g

\DeltaT = 8.5oC

C = Q/(w*\DeltaT)

C = 32.47J/(10 g* 8.5oC)

C = 0.382 J/goC

The specific heat of of the unknown metal is 0.382 J/goC.

Specific heat is the intensive property, from the standard specific heat capacity table, for Copper metal the specific heat capacity is 0.385 J/goC. So, the metal should be Copper.